# Interacting phantom dark energy: new accelerating scaling attractors

### v1

@@ III.2.1 Model I: $Q=\alpha H\rho_{m}$
Since the above autonomous system is invariant under the transformation $y\longrightarrow-y$, the phase space domain $\mathbb{D}$ is defined as $\mathbb{D}=\left\{(x,y,\lambda)\in\mathbb{R}^{3}:0\leq-x^{2}+y^{2}\leq 1,y\geq 0,-|\eta|\leq\lambda\leq|\eta|\right\}$.
@@ III.3.1 Model I: $Q=\alpha H\rho_{m}$
The above system is invariant under $y\longrightarrow-y$, and its phase space region $\mathbb{D}$ is defined by $\mathbb{D}=\left\{(x,y,\lambda)\in\mathbb{R}^{3}:0\leq-x^{2}+y^{2}\leq 1,y\geq 0\right\}$.
@@
  • The curve of critical points $F_{3}$ has $\lambda=0$, and it was absent for the exponential and hyperbolic potentials. In general these points correspond to scaling accelerating solutions, while for $x_{c}=\pm\frac{1}{\sqrt{2}}$ they correspond to completely matter-dominated points. Since, one eigenvalue is zero and the other two are negative, they are normally hyperbolic critical points. By applying the center manifold theorem Leon and Saridakis (2010) we deduce that these critical points behave as stable points for $\mu x_{c}>0$. Thus, these points can alleviate the coincidence problem for the model parameter $\alpha=-3$.
@@ III.3.2 Model II: $Q=\Gamma\rho_{m}$
where $\gamma=\Gamma/H_{0}$. The above autonomous system is invariant under $y\longrightarrow-y$, and the phase space domain is defined as $\mathbb{D}=\left\{(x,y,\lambda,z)\in\mathbb{R}^{4}:0\leq-x^{2}+y^{2}\leq 1,y\geq 0,0\leq z\leq 1\right\}$. The properties and features of the critical points are summarized in Table 6. They are the following.
@@
  • The curves of critical points $G_{3}$ exist for $|x_{c}|\leq\frac{1}{\sqrt{2}}$ with $\gamma<0$. They are normally hyperbolic, and by applying the center manifold theorem in order to determine their stability Leon and Saridakis (2010) we deduce that $G_{3}$ are stable for $\mu x_{c}>0$. Note that in general, for $|x_{c}|<\frac{1}{\sqrt{2}}$, these points have $10<\Omega_{\phi}<$ 1, and thus they correspond to a late-time stable accelerating solution with $\Omega_{m}/\Omega_{\phi}\neq 0$, namely with a scaling nature that can alleviate the coincidence problem. Finally, in the limiting case where $x_{c}=\pm\frac{1}{\sqrt{2}}$ they lead to $\Omega_{\phi}=0$ and thus to $\Omega_{m}=1$, corresponding to matter-dominated solutions.
@@ III.4.1 Model I: $Q=\alpha H\rho_{m}$
  • Point $I_{2}$ in its stability region has $0\leq\Omega_{\phi}\leq 1$, and thus $0\leq\Omega_{m}\leq 1$. Hence, since $\Omega_{\phi}$ and $\Omega_{m}$ can be of the same order, this point can alleviate the coincidence problem Amendola (2000); Chimento et al. (2003); Cai and Wang (2005); Pavon and Zimdahl (2005); Huey and Wandelt (2006); Hu and Ling (2006); del Campo et al. (2008, 2009). Moreover, for $\alpha=\frac{-13-\sqrt{88}}{3}$ or $\alpha=\frac{-33+\sqrt{88}}{3}$, it has $\Omega_{m}=1$, and in this case it corresponds to a matter-dominated solution.
@@ IV Summary and Conclusions
Our analysis extracted saddle matter-dominated points, stable dark-energy dominated points, and scaling accelerating solutions, that can attract the Universe at late times. As we discussed, some of the accelerating scaling attractors obtained for the last two potentials, when the local interaction is considered, are totally new. For completeness, in Fig. 4 we present the summary diagram that provides the qualitative behavior of the examined interacting phantom scenarios for different potentials.

### v2

@@ III.2.1 Model I: $Q=\alpha H\rho_{m}$
Since the above autonomous system is invariant under the transformation $y\longrightarrow-y$, the phase space domain $\mathbb{D}$ is defined as $\mathbb{D}=\left\{(x,y,\lambda)\in\mathbb{R}^{3}:0\leq-x^{2}+y^{2}\leq 1,y\geq 0,-|\eta|\leq\lambda\leq|\eta|\right\}$. The presence of $x$ in the denominator of equation (34) introduces a singularity at $x=0$ plane, causing the system to become ill-defined at this plane, and this can be resolved in a similar fashion as discussed in the previous subsections.
@@
Finally, investigating the qualitative behaviour at the critical points at infinity (see Appendix B.1) reveals that the phase space domain $\mathbb{D}$ contains four such critical points, namely $C_{1}^{\pm\infty}\left(\pm\frac{1}{\sqrt{2}},\frac{1}{\sqrt{2}},0,0\right)$ and $C_{2}^{\pm\infty}\left(\pm\sqrt{\frac{3}{8}},\sqrt{\frac{3}{8}},\mp\frac{1}{2},0\right)$. The first two points are unstable, while the remaining two exhibit saddle behavior. At these points, we deduce that $q\longrightarrow-\infty$.
@@ III.2.2 Model II: $Q=\Gamma\rho_{m}$
Since $x$ and $(1-z)$ appear in the denominator of the evolution equation (37), the system exhibits singularities at $x=0$ and $z=1$ surfaces, making it ill-defined along these surfaces.
@@
Finally, with the help of the the Poincaré compactification it follows that the system encounters two critical points at infinity, namely $D^{+\infty}\Bigl{\{}\left(y_{1},y_{2},y_{3},y_{4},y_{5}\right)\in\mathbb{S}^{4}:y_{5}=0,y_{4}=0,y_{1}^{2}=y_{2}^{2},y_{1}>0,y_{2}>0\Bigr{\}}$, and $D^{-\infty}\Bigl{\{}\left(y_{1},y_{2},y_{3},y_{4},y_{5}\right)\in\mathbb{S}^{4}:y_{5}=0,y_{4}=0,y_{1}^{2}=y_{2}^{2},y_{1}<0,y_{2}>0\Bigr{\}}$, where $q\longrightarrow-\infty$. The analysis reveals that these critical points exhibit unstable behavior. The details of the analysis are presented in Appendix B.2.
@@ III.3.1 Model I: $Q=\alpha H\rho_{m}$
The above system is invariant under $y\longrightarrow-y$, and its phase space region $\mathbb{D}$ is defined by $\mathbb{D}=\left\{(x,y,\lambda)\in\mathbb{R}^{3}:0\leq-x^{2}+y^{2}\leq 1,y\geq 0\right\}$. The explicit presence of $x$ in the denominator of equation (42) leads to a singularity at $x=0$ plane, where the system is no longer well-defined.
@@
  • The curve of critical points $F_{3}$ has $\lambda=0$, and it was absent for the exponential and hyperbolic potentials. In general these points correspond to scaling accelerating solutions, while for $x_{c}=\pm\frac{1}{\sqrt{2}}$ they correspond to completely matter-dominated points. Since, one eigenvalue is zero and the other two are negative, they are normally hyperbolic critical points. By applying the center manifold theorem [103] we deduce that these critical points behave as stable points for $\mu x_{c}>0$. Thus, these points can alleviate the coincidence problem for the model parameter $\alpha=-3$.
Finally, from the critical point analysis at infinity (see Appendix C.1), we deduce that the phase space domain $\mathbb{D}$ includes four such critical points, namely $E_{1}^{\pm\infty}\left(\pm\frac{1}{\sqrt{2}},\frac{1}{\sqrt{2}},0,0\right)$, and $E_{2}^{\pm\infty}\left(\pm\sqrt{\frac{2}{7}},\sqrt{\frac{2}{7}},\mp\sqrt{\frac{3}{7}},0\right)$. The first two points are unstable, while the last two are saddles. At these critical points, it follows that $q\longrightarrow-\infty$.
@@ III.3.2 Model II: $Q=\Gamma\rho_{m}$
where $\gamma=\Gamma/H_{0}$. The above autonomous system is invariant under $y\longrightarrow-y$, and the phase space domain is defined as $\mathbb{D}=\left\{(x,y,\lambda,z)\in\mathbb{R}^{4}:0\leq-x^{2}+y^{2}\leq 1,y\geq 0,0\leq z\leq 1\right\}$. The dynamical system has two singular surfaces, namely $x=0$ and $z=1$, in which it is ill-defined. The properties and features of the critical points are summarized in Table 6. They are the following.
@@
  • The curves of critical points $G_{3}$ exist for $|x_{c}|\leq\frac{1}{\sqrt{2}}$ with $\gamma<0$. They are normally hyperbolic, and by applying the center manifold theorem in order to determine their stability [103] we deduce that $G_{3}$ are stable for $\mu x_{c}>0$. Note that in general, for $|x_{c}|<\frac{1}{\sqrt{2}}$, these points have $10<\Omega_{\phi}<$ 1, and thus they correspond to a late-time stable accelerating solution with $\Omega_{m}/\Omega_{\phi}\neq 0$, namely with a scaling nature that can alleviate the coincidence problem. Finally, in the limiting case where $x_{c}=\pm\frac{1}{\sqrt{2}}$ they lead to $\Omega_{\phi}=0$ and thus to $\Omega_{m}=1$, corresponding to matter-dominated solutions.
Lastly, the system encounters two critical points at infinity, namely $F^{+\infty}\Bigl{\{}\left(y_{1},y_{2},y_{3},y_{4},y_{5}\right)\in\mathbb{S}^{4}:y_{5}=0,y_{4}=0,y_{1}^{2}=y_{2}^{2},y_{1}>0,y_{2}>0\Bigr{\}}$ and $F^{-\infty}\Bigl{\{}\left(y_{1},y_{2},y_{3},y_{4},y_{5}\right)\in\mathbb{S}^{4}:y_{5}=0,y_{4}=0,y_{1}^{2}=y_{2}^{2},y_{1}<0,y_{2}>0\Bigr{\}}$, where $q\longrightarrow-\infty$. These critical points exhibit unstable behavior. The details of the analysis are presented in Appendix C.2.
@@ III.4.1 Model I: $Q=\alpha H\rho_{m}$
The system exhibits a singular surface at $x=0$, where it becomes ill-defined. The critical points, the existence and stability conditions as well as the corresponding values of $\Omega_{\phi}$ and $w_{tot}$ are summarized in Table 7, while the corresponding eigenvalues are displayed in Table 8. For this scenario we have the following critical points.
@@
  • Point $I_{2}$ in its stability region has $0\leq\Omega_{\phi}\leq 1$, and thus $0\leq\Omega_{m}\leq 1$. Hence, since $\Omega_{\phi}$ and $\Omega_{m}$ can be of the same order, this point can alleviate the coincidence problem [6, 7, 8, 9, 10, 11, 12, 13]. Moreover, for $\alpha=\frac{-13-\sqrt{88}}{3}$ or $\alpha=\frac{-33+\sqrt{88}}{3}$, it has $\Omega_{m}=1$, and in this case it corresponds to a matter-dominated solution.
@@
Finally, we observe that the system possesses two critical points at infinity, namely $G^{\pm\infty}\Bigl{\{}\left(y_{1},y_{2},y_{3},y_{4},y_{5}\right)\in\mathbb{S}^{4}:y_{5}=0,y_{4}>0,y_{3}>0,y_{2}>0,y_{1}=\pm\frac{1}{\sqrt{2}}\Bigr{\}}$, exhibiting unstable behavior and having $q\longrightarrow-\infty$. The details of the analysis are presented in Appendix D.1.
@@ III.4.2 Model II: $Q=\Gamma\rho_{m}$
There are two singular surfaces in the dynamical system, at $x=0$ and $z=1$ respectively, at which the system is ill-defined.
@@
Lastly, the Poincaré compactification of the system results in two critical points at infinity, given by, $J^{\pm\infty}\Bigl{\{}\left(y_{1},y_{2},y_{3},y_{4},y_{5},y_{6}\right)\in\mathbb{S}^{5}:y_{6}=0,y_{5}=0,y_{4}>0,y_{3}>0,y_{2}>0,y_{1}=\pm\frac{1}{\sqrt{2}}\Bigr{\}}$, which exhibit unstable behavior, and having $q\longrightarrow-\infty$. The details are presented in Appendix D.2.
@@ IV Summary and Conclusions
Our analysis extracted saddle matter-dominated points, stable dark-energy dominated points, and scaling accelerating solutions, that can attract the Universe at late times. As we discussed, some of the accelerating scaling attractors obtained for the last two potentials, when the local interaction is considered, are totally new. Furthermore, performing a detailed analysis at infinity it has been shown that there are no stable critical points at infinity, and thus they cannot attract the Universe at late times. For completeness, in Fig. 4 we present the summary diagram that provides the qualitative behavior of the examined interacting phantom scenarios for different potentials.